OpenStudy (anonymous):

There are two girls. Their ages happen to be consecutive odd integers. Girl one is the younger of the two. Forty less than five times girl one's age is equal to three time the age of girl two. How old is girl two?

OpenStudy (anonymous):

x girl 1 x+2 girl 2 (x)5-40=3(x+2) 5x-40=3x+2 2x=42 x=21

OpenStudy (anonymous):

girl 2 = 23

OpenStudy (anonymous):

Let g1=2n+1, g2=2n+3, 5g1-40=3g2. Solve the three simultaneous equations for g1, g2 and n. {g1 = 23, g2 = 25, n = 11}

OpenStudy (anonymous):

What does the 2n stand for?

OpenStudy (anonymous):

The expression 2n+1 where n = 1,2,3, ... will always produce an odd number. The problem ages were two "consecutive odd numbers". n is a solved number such that when plugged into the expressions 2n+1 and 2n+3, will yield the ages of g1 and g2 respectively.

OpenStudy (anonymous):

So, how would the equation look? Because i just tried the 5g1-40=3g2 and i didnt get the right number. Im not really sure how to set it up right. The math should be easy for me just setting it up is stumping me.

OpenStudy (anonymous):

By the way, I don't think shinigami1m's solution is correct for g2.

OpenStudy (anonymous):

Well im not sure, i should have mentioned that this was a multiple choice question, and there are 4 choices which are 1) 20 2) 21 3) 23 4) 25. I jumped at the answer because it thought it made sense but i double checked it and im not sure because i dont see where the first consec. odd integer comes in there. I just see 5x.

OpenStudy (anonymous):

5*23 - 40 = 3*25 ? 115 - 40 = 75 ? 75 = 75 Yes.

OpenStudy (anonymous):

Made an error in my first posting of the following: 5 g1 - 40 = 3 g2 5 (2 n + 1) - 40 = 3 (2 n + 3) n = 11

OpenStudy (anonymous):

siety wrong algebra x girl 1 x+2 girl 2 (x)5-40=3(x+2) 5x-40=3x+6 2x=46 x=23 so the 2nd girl is 25 years old

OpenStudy (anonymous):

Okay i went through this and i finally arrived at 25. Thanks to both of you for helping me! :)